OpenStudy (anonymous):

Serena parachutes from a plane at a height of 20,000 ft above sea level. Several minutes later she lands on the ground 15 ft below sea level . What was the total vertical distance that Serena parachuted?

OpenStudy (anonymous):

Will give medal to who ever gets it corrected

OpenStudy (anonymous):

DTotal=D1+D2 D1=20000 D2=15

OpenStudy (anonymous):

so is it 20,015?

OpenStudy (anonymous):

20000 - (-15) = 20000 + 15 = 20015

OpenStudy (anonymous):

Yes Correct!

OpenStudy (anonymous):

i have a few more stick around

OpenStudy (anonymous):

The variable t represents the temperature at 3 p.m. Which equation would best model this word problem? The temperature at the summit of Mt Mitchell dropped 23°F between 3 p.m. and 9 p.m. If the temperature at 9 p.m. was –10°F, what was the temperature at 3 p.m.?

OpenStudy (anonymous):

can u guys see that?

OpenStudy (anonymous):

t - 23 = -10 t = -10 + 23 t = 13